### v1

# RGB-Event based Pedestrian Attribute Recognition: A Benchmark Dataset and An Asymmetric RWKV Fusion Framework

Abstract
Existing pedestrian attribute recognition methods are generally developed based on RGB frame cameras. However, these approaches are constrained by the limitations of RGB cameras, such as sensitivity to lighting conditions and motion blur, which hinder their performance. Furthermore, current attribute recognition primarily focuses on analyzing pedestrians’ external appearance and clothing, lacking an exploration of emotional dimensions. In this paper, we revisit these issues and propose a novel multi-modal RGB-Event attribute recognition task by drawing inspiration from the advantages of event cameras in low-light, high-speed, and low-power consumption. Specifically, we introduce the first large-scale multi-modal pedestrian attribute recognition dataset, termed EventPAR, comprising 100K paired RGB-Event samples that cover 50 attributes related to both appearance and six human emotions, diverse scenes, and various seasons. By retraining and evaluating mainstream PAR models on this dataset, we establish a comprehensive benchmark and provide a solid foundation for future research in terms of data and algorithmic baselines. In addition, we propose a novel RWKV-based multi-modal pedestrian attribute recognition framework, featuring an RWKV visual encoder and an asymmetric RWKV fusion module. Extensive experiments are conducted on our proposed dataset as well as two simulated datasets (MARS-Attribute and DukeMTMC-VID-Attribute), achieving state-of-the-art results. The source code and dataset will be released on https://github.com/Event-AHU/OpenPAR
1 Introduction
Existing Pedestrian Attribute Recognition (PAR) [44] systems are typically developed using RGB frame cameras, with the goal of enabling a model to accurately select matching attributes from a predefined list based on a given pedestrian image or video. Common attributes include appearance information such as hairstyle, gender, upper and lower body clothing, and carried objects. With the advancement of artificial intelligence technology, pedestrian attribute recognition has garnered increasing attention and is now widely applied in various human-centric visual tasks, including pedestrian detection and tracking [43, 48, 27, 28], person re-identification [42, 13, 55], and text-image retrieval [10, 53, 64] for pedestrians. Despite significant progress, the overall performance of these systems remains constrained by the inherent limitations of RGB cameras, such as sensitivity to lighting conditions, clutter background, and susceptibility to motion blur. Interestingly, while multi-modal fusion has been extensively adopted in other computer vision tasks [52, 50, 35], it is surprising that pedestrian attribute recognition still predominantly relies on single-modality RGB cameras. This limitation highlights an opportunity for innovation through the integration of additional modalities to address the existing performance bottlenecks.
Figure 1: (left) Comparison between existing PAR datasets and our newly proposed EventPAR; (right) Imaging principle and results of Frame and Event cameras.
On the other hand, the pedestrian attributes defined in existing PAR benchmark datasets focus on describing the visible appearance cues, but none of them consider the invisible emotion information, which is also very important for human-centric perception. Recognizing emotions such as anxiety, anger, or joy can help the system better understand pedestrian behavioral intent and enhance the naturalness of human-computer interaction, improve safety and risk assessment, and support personalized services. Incorporating emotion analysis into pedestrian attribute recognition tasks can make the system smarter and more human-centric, while also providing support for a broader range of practical application scenarios.
In this work, we first propose to recognize human attributes using a DVS346 event camera that outputs aligned frame-event streams. This is because event cameras have inherent advantages in high dynamic range, high temporal resolution, and low power consumption. However, event streams contain some noise and are less sensitive to static objects. Therefore, combining them with RGB frame cameras is necessary to achieve more robust and accurate results. Due to the lack of datasets in this area, we first propose a large-scale multi-modal RGB-Event pedestrian attribute recognition benchmark dataset, termed EventPAR. This dataset covers common human appearance attributes as well as six human emotional attributes, totaling 50 attributes and 100K paired RGB-Event samples, as shown in Fig. 1. These videos were collected under various challenging scenarios, including different lighting conditions, motion blur, object occlusion, and adverse weather conditions. In addition to natural challenging factors, we also artificially introduced some adversarial attack noises and degradation strategies, such as various types of noise and occlusion. To provide a solid benchmark for future research, we re-train and report 18 mainstream PAR algorithms, as shown in Table 3.
Based on the proposed EventPAR dataset, we introduce a novel multi-modal RGB-Event pedestrian attribute recognition framework leveraging the RWKV model [8], which effectively integrates event streams to further enhance pedestrian attribute recognition results from the traditional RGB frame. Specifically, given the input RGB image and event streams, we first employ an RWKV encoder to extract visual features from both modalities. These features are then fed into an asymmetric RWKV fusion module, designed to better fuse spatial features from RGB frames with temporal information from event streams. This fusion module efficiently removes redundant event tokens while achieving interactive fusion with RGB features. Finally, a linear layer maps the fused features to multi-label attribute predictions. We extensively validate the proposed framework on three datasets and achieve state-of-the-art performance.
To sum up, we draw the main contributions of this paper as the following three aspects:
1). We propose to recognize the pedestrian attributes using an event camera and construct a large-scale benchmark dataset termed EventPAR. It contains 100K aligned RGB frame and event streams, collected from multiple scenarios, illuminations, etc.
2). We propose a novel RWKV-based framework for the frame-event fusion based attribute recognition. A novel asymmetric RWKV fusion module fuses the RGB frame and event streams effectively and efficiently.
3). We build a large-scale benchmark for the EventPAR dataset by re-training 18 representative and strong PAR algorithms. This benchmark will be a good platform to boost the development of event-based attribute recognition further.
In addition, we also conduct extensive experiments on three PAR benchmark datasets, including the newly proposed EventPAR, MARS-Attribute and DukeMTMC-VID-Attribute dataset. More experiments fully validated the effectiveness of our proposed model for the frame-event pedestrian attribute recognition.
2 Related Works
2.1 Pedestrian Attribute Recognition
Pedestrian Attribute Recognition aims to identify and classify multiple attributes of pedestrians, such as age, gender, clothing type, and accessories, from images or videos. For image-based PAR tasks, CLEAR [1] combines pre-trained language models to generate pseudo-descriptions for attribute queries and extracts robust features using Cross Transformers and adapters. HPNet [29] and DAHAR [49] focus on attribute-related regions via attention mechanisms. PARformer [9] extracts features by combining global and local perspectives. VTB [5] introduces a text encoder for pedestrian attribute recognition tasks. Visual Prompt Tuning (VPT) [16] and CLIP-Adapter [11] enhance few-shot recognition performance by fine-tuning frozen backbones and combining original CLIP knowledge with limited-shot knowledge, respectively.
Video-based PAR, compared to image-based methods, is a relatively new research topic. Researchers have introduced multi-task models, temporal pooling, and sparse temporal attention modules, among other techniques. Chen et al. [3] proposed a multi-task model with an attention module to handle each attribute per frame. Specker et al. [38] introduced global features to incorporate information from different frames. Li et al. [21] proposed a sparse temporal attention module to select unobstructed frames, enhancing robustness. Thakare et al. [40] calculated the mutual correlations of attribute predictions from pedestrian images across different views, while Liu et al. [30] introduced a spatio-temporal saliency module to capture correlations between attributes in both spatial and temporal domains. Different from these works, this paper firstly exploits the multi-modal PAR task using the event camera and proposes a large-scale benchmark dataset and RWKV-based framework for this task.
2.2 Event-based Vision
Event cameras have been widely used in computer vision and multi-modal related tasks. Specifically, in the image restoration and enhancement tasks, EventSR [41] proposes a novel end-to-end super-resolution image reconstruction framework that reconstructs low-resolution images from event streams, enhances image quality, and upsamples the enhanced images. EG-VSR [31] learns spatial-temporal coordinates and implicit neural representations from RGB frames and event features, leveraging the high temporal resolution of events to achieve video super-resolution at arbitrary resolutions. Jiang et al. [18] and colleagues have introduced a method that effectively restores motion-blurred video sequences from event-based camera data, achieving this by integrating global and local scale visual and temporal knowledge, along with a differentiable directional event filtering module. Jiang et al. [17] proposed a posterior attention module that adjusts standard attention using the prior knowledge provided by event data, which improves the segmentation performance for moving objects. Inspired by these works, in this paper, we attempt to achieve high-performance PAR by fusing the event stream and RGB frames.
2.3 RWKV Model
The Receptance Weighted Key Value (RWKV) [33] model was first proposed in the field of natural language processing (NLP). Before the emergence of RWKV, Transformers were extensively used in NLP. However, Transformers suffer from quadratic time complexity with respect to sequence length, which significantly limits their efficiency, particularly when processing long sequences. Inspired by the linear attention mechanism in the Attention Free Transformer (AFT) [37], RWKV introduced the WKV attention mechanism. At the same time, the token shift operation was proposed to capture local context. This combination allows RWKV to take advantage of both the parallel processing attention mechanism of Transformers and the sequential time-series pattern of Recurrent Neural Networks (RNNs). Vision-RWKV [58] then extended the RWKV model to the computer vision field. BSBP-RWKV [51] proposes an efficient medical image segmentation framework that integrates Perona-Malik Diffusion for noise suppression with boundary preservation and an RWKV-based architecture to reduce computational complexity. RWKV-SAM [56] proposes an efficient segment-anything framework with a mixed convolution-RWKV backbone and multiscale token decoder, using a joint training benchmark to achieve twice the inference speed and superior segmentation quality compared to Transformer and vision Mamba baselines. Different from these works, this paper first adopts the RWKV model for the pedestrian attribute recognition task.
Figure 2: An illustration of representative samples in our newly proposed EventPAR dataset. The left side displays RGB and Event bimodal data collected during summer and winter seasons, as well as the synthetic challenges incorporated into the dataset. The right side presents examples of the six emotion-related attributes.
3 EventPAR Benchmark Dataset
3.1 Protocols
In this work, we follow the guidelines below when constructing the EventPAR benchmark dataset: 1). Multi-Modal: Unlike existing frame-based PAR datasets, we utilize the DVS346 camera for data collection, obtaining spatiotemporal aligned multi-modal data comprising visible light and event streams. It is the first multi-modal PAR database, laying a solid data foundation for subsequent research. 2). Emotion Attributes: Unlike conventional PAR datasets that primarily focus on pedestrians’ appearance attributes, our new dataset incorporates six fundamental human emotional attributes, i.e., Happiness, Sadness, Anger, Surprise, Fear, and Disgust. This addition effectively fills the gap in the attribute recognition domain by addressing the lack of discussion on the emotional analysis dimension. 3). Diverse Scenes and Cross-Seasonal: Different from existing PAR datasets that involve short-term data collection, our data acquisition spanned several months, encompassing data from different seasons (summer and winter). Additionally, it covers a variety of scenes and weather conditions, such as daytime and nighttime scenes, as well as weather conditions like sunny and rainy days. These elements significantly enhance the diversity and challenge of the dataset, providing better validation of a model’s fitting ability and generalization capability. 4). Large Scale: It contains 100K pedestrian samples, 12 attribute groups, and 50 fine-grained pedestrian attributes. It is one of the largest PAR datasets, comparable to PA100K in scale. 5). Simulated Complex Real-world Environments: Our dataset is carefully designed to incorporate a wide range of challenges, such as variations in illumination, motion blur, object occlusion, and adverse weather conditions, all of which simulate the complex real-world difficulties encountered in pedestrian attribute recognition. To further increase the complexity and practical relevance of the dataset, we employ adversarial attack techniques [2, 32, 59] during data processing, adding an extra layer of difficulty. This approach not only enhances the robustness of pedestrian attribute recognition technologies but also ensures their applicability in real-world scenarios, as shown in Fig. 2.
Table 1: Comparison with existing PAR benchmark datasets. ##\##Att. is short for the number of Attributes. Deg and Adv is short for Degradation and Adversary, respectively.
Dataset Year ##\##Att. ##\##Samples Video Frame Event Emotion Deg Adv
PETA [7] 2014 61 19,000
WIDER [26] 2016 14 57,524
RAPv1 [23] 2016 69 41,585
PA100K [29] 2017 26 100,000
RAPv2 [24] 2019 76 84,928
MSP60K [20] 2024 57 60,015
MARS-Attribute [61] 2016 43 16,360
DukeMTMC-VID [34] 2017 36 4,832
EventPAR (Ours) 2025 50 100,000
Table 2: Attribute groups and details defined in our proposed EventPAR dataset.
Attribute Group Details
#01 Gender Male, Female
#02 Age Child, Adult, Elderly
#03 Body Size Fat, Normal, Thin
#04 Viewpoint Front, Back, Side
#05 Head Long Hair, Black Hair, Hat, Glasses, Mask, Scarf, Headphones
#06 Upper Body Short Sleeves, Long Sleeves, Shirt, Coat, Cotton-padded Coat
#07 Lower Body Trousers, Shorts, Jeans, Long Skirt, Short Skirt, Casual Pants, Dress
#08 Shoes Casual Shoes, Other Shoes
#09 Accessory Backpack, Shoulder Bag, Hand Bag, Umbrella, Cup, Cellphone
#10 Posture Walking, Running, Standing, Sitting
#11 Activity Smoking, Reading
#12 Emotion Happiness, Sadness, Anger, Surprise, Fear, Disgust
3.2 Data Collection and Annotation
Based on the aforementioned protocols, we collect the EventPAR dataset using a DVS346 event camera, which outputs both RGB frames and event streams. The entire data collection process spanned several months, covering multiple seasons from summer to winter. Approximately 20 students participated in the data collection and subsequent annotation process. As shown in Table 2, our dataset contains 100K spatio-temporal aligned RGB frames and event streams. Each sample in our dataset is annotated with 50 attributes and can be divided into 12 attribute groups, including Gender, Age, Body Size, Viewpoint, Head, Upper Body, Lower Body, Shoes, Accessory, Posture, Activity, and Emotion. More details can be found in Table 2.
By following the MSP60K dataset [20], we also introduced the data quality degradation strategy to evaluate the robustness of models against challenging factors, including changes in lighting, random occlusions, blurring, noise, and adversarial disturbance. Our dataset also exhibits a long-tail distribution, as depicted in Fig. 3 (a), which aligns with the attribute distribution typically observed in real-world data. It is evident that attributes such as Adult, Back Hair, and Normal Size are among the top-3 most frequent, while attributes like Smoking, Child, and Elderly rank at the bottom. We also provide the co-occurrence matrix of our EventPAR dataset as shown in Fig. 3 (b).
Figure 3: (a) Attributes Distribution: Bar graph showing the prevalence of individual attributes across the dataset; (b) Co-occurrence Matrix of Attributes: Logarithmic heatmap showing the co-occurrence frequency of attribute pairs.
3.3 Benchmark Baselines
Based on our newly proposed dataset, we retrained and tested 18 mainstream PAR models, providing a solid baseline for future work on the EventPAR dataset. Specifically, the following algorithms were included: 1). CNN-based: DeepMAR [22],Zhou et al. [62], RethinkPAR [15], SSCNet [14], SSPNet [36]. 2). Transformer-based: DFDT [60], PARFormer [9]. 3). Mamba-based: MambaPAR [47], MaHDFT [46]. 4). Human-Centric Pre-Training Models for PAR: PLIP [65], HAP [57]. 5). Visual-Language Models for PAR: VTB [4], Label2Label [25], PromptPAR [45], SequencePAR [19].
Figure 4: An overview of our proposed OTN-RWKV guided RGB-Event fusion for pedestrian attribute recognition.
4 Methodology
4.1 Overview
To enhance recognition accuracy and robustness, in this paper, we propose a video-based multi-modal pedestrian attribute recognition framework, as illustrated in Fig. 4. Our method leverages event data as the auxiliary modality to compensate for information loss in the RGB modality under challenging environments, thereby improving the stability of recognition. With the powerful modeling capabilities of Vision-RWKV, we propose an OTN-RWKV guided multimodal fusion framework. Concretely, we first employ the Visible-RWKV encoder and Event-RWKV encoder to extract the specific features from RGB and event data, respectively. The extracted features are subsequently processed by the proposed OTN-RWKV fusion module to facilitate the interaction and integration of multimodal information. The resulting fused representation is then passed to a classifier for accurate attribute prediction. Please refer to the following for more detailed information.
4.2 Network Architecture
$\bullet$ Input Encoding : In this work, event modality is used to assist the RGB video in overcoming information loss in complex environments. Given a sequence of an RGB video denote as $X_{r}\in\mathbb{R}^{T\times H\times W\times C}$, where $T$ is the total number of frames, and $H$, $W$, $C$ correspond to the height, width, and the number of channels of each frame, respectively. We take event streams as =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$, each event point $e_{j}$ can be denoted as $\{x,y,t,p\}$, representing spatial coordinates, timestamps, and polarity, respectively. To align with the RGB sequence, we stack the event streams into event frames that are temporally synchronized with the RGB frames based on exposure times. Therefore, we can obtain the event frame sequence $X_{e}\in\mathbb{R}^{T\times H\times W\times C}$.
By inputting a sequence of RGB frames $X_{r}\in\mathbb{R}^{T\times H\times W\times C}$ and a corresponding sequence of event frames $X_{e}\in\mathbb{R}^{T\times H\times W\times C}$, we first apply a partitioning operation to process data for both modalities, transforming each frame into ${HW}/p^{2}$ patches, where $p$ denotes the patch size. Then, the patches are added with Position Embedding (P.E.), which are used to encodes spatial information to obtain RGB tokens $X_{r}\in\mathbb{R}^{T\times N\times C}$ and event tokens $X_{e}\in\mathbb{R}^{T\times N\times C}$ separately, where $N$ is the number of tokens.
$\bullet$ RWKV Encoder: In this paper, we employ Vision-RWKV [8] as our visual encoder, which adopts a block-stacked design with $L$ layers to enhance its encoding capability, where each block consists of a Spatial-Mix module and a Channel-Mix module. In each layer, RGB tokens $X_{r}\in\mathbb{R}^{T\times N\times C}$ and event visual tokens $X_{e}\in\mathbb{R}^{T\times N\times C}$ are first fed into the Spatial-Mix module, which performs linear complexity global attention computation. Specifically, as shown in the top right of Fig. 4, the input visual tokens first undergo a token Shift operation to capture local context information before being fed into three parallel linear layers to obtain the matrices $R_{s}$,$K_{s}$,$V_{s}\in\mathbb{R}^{T*N\times C}$. The formulas are as follows,
$R_{\text{s}}$ $={\rm Q\mbox{-}Shift}_{R}(X_{i})W_{R},$ (1)
$K_{\text{s}}$ $={\rm Q\mbox{-}Shift}_{K}(X_{i})W_{K},$
$V_{\text{s}}$ $={\rm Q\mbox{-}Shift}_{V}(X_{i})W_{V},\quad i\in(r,e)$
where the Q-Shift function is a specialized token shift mechanism specifically designed for vision tasks. After that, $K_{s}$ and $V_{s}$ are used to calculate the global attention result $wkv\in\mathbb{R}^{T*N\times C}$ by the Bi-WKV module, which is a linear complexity bidirectional attention mechanism. Subsequently, the $wkv$ is multiplied with $\sigma(R)$, which controls the output $O_{i}$ probability,
$O_{i}$ $=(\sigma(R_{s})\odot whv)W_{O},$ (2)
[…]
where the operator $\sigma$ represents the sigmoid function, while $\odot$ denotes element-wise multiplication. After the output linear projection, the features are subjected to layer normalization to enhance stability and facilitate effective training.
Subsequently, the tokens are fed into the Channel-Mix module to achieve channel-wise fusion, as shown in the bottom right of Fig. 4. The parameters $R_{c}$ and $K_{c}$ are computed similarly to the Spatial-Mix before the final output,
[…]
the gating mechanism $\sigma(R_{c})$ is used to modulate the output $O^{\prime}_{i}$. To maintain information integrity and promote gradient flow, residual connections are introduced to fuse the module outputs, thereby effectively alleviating the vanishing gradient problem. This process can be expressed as,
$O_{i}^{\prime}$ $=(\sigma(R_{c})\odot V_{c})W_{O},$ (4)
$V_{c}$ $=\text{SquaredReLU}(K_{c})W_{V},$
Figure 5: Illustration of our proposed OTN-RWKV for RGB-Event fusion.
$\bullet$ OTN-RWKV: Due to the asynchronous nature of event cameras, the event streams tend to be highly dense, resulting in stacked event frames containing a significant amount of redundant information. This redundancy not only increases computational overhead but also causes information ambiguity, thereby diminishing the effectiveness of event data as auxiliary information. To address this issue, we propose the OTN-RWKV fusion module, which first applies the similarity-based filtering on the visual tokens output $O_{e}^{\prime}$ of event modality, removing highly redundant tokens and achieving one-to-N multimodal fusion. Specifically, a similarity matrix is employed to identify the top-$K$ most similar token pairs, from which only the most representative tokens are retained. This approach enhances data efficiency and improves the quality of multimodal fusion. The implementation process is as follows,
$O_{e}^{{}^{\prime\prime}}$ $=\text{KNP}_{filter}(\text{sim}(O_{e}^{\prime},O_{e}^{\prime}))\odot O_{e}^{\prime},$ (5)
where $O_{e}^{{}^{\prime\prime}}\in\mathbb{R}^{N^{\prime}\times C}$ represents the tokens for the informative filtered event. After filtering, both modalities are aligned to have the same number of tokens, facilitating effective multimodal fusion.
Although traditional fusion methods (e.g., concatenation, addition, and 1×1 convolution) generally exhibit strong compatibility and generalization across different models, they frequently result in information loss, limiting their effectiveness in preserving fine-grained multimodal details. Inspired by the concept of cross attention and LCR[54], in this work, we further propose a novel interactive fusion strategy. As shown in Fig. 5, we first apply the Q-shift operation separately to the two types of visual tokens, corresponding to Eq.1 and Eq. 3, respectively. Subsequently, $V_{s}$, $R_{c}$, and $K_{c}$ are used to compute the cross-attention results through a linear complexity bidirectional attention mechanism,
$O_{fusion}=\sigma(R_{s})\odot\text{$LN$}(K_{s}\odot\text{Bi-}WKV)+O_{e}^{{}^{\prime}},$ (6)
[…]
where $t$ represents the $t$-th token of the input, $M$ represents the total number of tokens, $u$ is a C-dimensional learnable vector that acts as a bonus indicating the current token, and $LN$ refers to LayerNorm. Unlike LCR[54], which uses the learnable vector $w$, we replace $w$ with the output $V_{s}$ from Eq. 1 to better focus on the current sample.
4.3 Loss Function
After the effective fusion of the two modalities using the OTN-RWKV module, the resulting representation is fed into an average pooling layer and then a linear classification head, yielding the final attribute prediction $P_{attr}$. In the training phase, we adopt the weighted cross-entropy loss (WCE Loss)[4] to alleviate the distribution imbalance between pedestrian attributes, which is widely used by attribute recognition,
$\mathcal{L}_{wce}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\omega_{j}\left(y_{ij}\log(p_{ij})+(1-y_{ij})\log(1-p_{ij})\right),$ (8)
here, $N$ represents the total number of samples, $M$ denotes the number of attributes, and $\omega_{j}$ is the weight for the $j$-th attribute. $y_{ij}$ and $p_{ij}$ denote the ground truth and predicted result, respectively.
5 Experiments
5.1 Datasets & Evaluation Metric
In this study, we conduct a comprehensive benchmark of 18 pedestrian attribute recognition methods, representing the most important models in the field of pedestrian attribute recognition. Additionally, the performance of our method is compared with existing state-of-the-art (SOTA) pedestrian attribute recognition (PAR) methods across our benchmark and two publicly available datasets: MARS-Attribute Dataset [61] and DukeMTMC-VID-Attribute Dataset [34].
$\bullet$ MARS-Attribute Dataset is an annotated dataset for pedestrian attribute recognition, containing multiple multi-label and binary attributes such as pedestrian action, orientation, clothing color, gender, and others. These attributes are decomposed into 43 binary attributes to enhance the performance of pedestrian attribute recognition models for training and testing. The dataset is divided into a training subset and a testing subset, with 8,298 and 8,062 tracklets, respectively. Each tracklet contains approximately 60 frames on average. The dataset includes 625 and 626 distinct pedestrian identities for training and testing.
$\bullet$ DukeMTMC-VID-Attribute Dataset is an extension of the DukeMTMC-VID dataset, specifically designed for pedestrian attribute recognition. This dataset includes various annotated pedestrian attributes aimed at improving pedestrian re-identification performance. It features 2,032 identities and 16,522 video sequences captured across multiple scenarios, with a focus on real-world and challenging environments. In addition, the dataset provides rich attribute annotations. By splitting the multi-label attributes into binary attributes, the dataset includes a total of 36 binary attributes for training and testing. The training subset contains 702 distinct pedestrian identities and 16,522 images, while the testing subset includes 17,661 images corresponding to 702 pedestrians. The DukeMTMC-VID-Attribute dataset is widely used for evaluating attribute-based pedestrian re-identification models, offering valuable insights into how attributes influence recognition accuracy under varying conditions.
For the evaluation of our and the compared PAR models, we use five widely adopted evaluation metrics to assess performance, including: mean Accuracy (mA), Accuracy (Acc), Precision (Prec), Recall, and F1-score (F1), which can be expressed as:
$\centering\small Accuracy=\dfrac{TP+TN}{FP+FN+TP+TN}\@add@centering$ (9)
$\centering\small Precision=\dfrac{TP}{FP+TP},~{}~{}Recall=\dfrac{TP}{TP+FN}\@add@centering$ (10)
$\centering\small F1=\dfrac{2\times Recall\times Precision}{Recall+Precision}\@add@centering$ (11)
where $TP$ denotes the number of correctly predicted positive samples, $TN$ is the number of correctly predicted negative samples, $FP$ and $FN$ denote the number of false positive and false negative samples, respectively.
Table 3: Comparison with public methods on our datasets. The first and second are shown in red and blue, Size indicates the size of checkpoint file, and ##\##MS means the GPU memory used for optimization, Test time refers to the time spent on the test set.
Methods Publish Code mA Acc Prec Recall F1 Test Time Params (MB) Flops (GB) Size (MB) ##\##MS (MB)
#01  DeepMAR [22] ACPR15 URL 71.15 72.39 86.42 78.93 82.51 46s 23 4 181 846
#02  ALM  [39] ICCV19 URL 71.31 68.80 78.82 75.97 77.37 21s 17 45 66 484
#03  Strong Baseline [15] arXiv20 URL 73.75 70.63 81.34 78.26 79.32 91s 24 4 91 788
#04  RethinkingPAR [15] arXiv20 URL 71.23 69.00 80.80 77.16 78.94 57s 24 4 91 105
#05  SSCNet [14] ICCV21 URL 72.12 69.95 78.36 80.10 79.22 24s 24 3 90 854
#06  VTB [4] TCSVT22 URL 76.96 73.81 80.74 83.15 81.67 441s 93 15 333 114
#07  Label2Label [25] ECCV22 URL 72.47 69.54 78.75 79.14 78.55 43s 66 12 829 3788
#08  DFDT [60] EAAI22 URL 70.14 70.95 78.71 81.82 80.24 193s 88 26 669 1884
#09  Zhou et al. [62] IJCAI23 URL 70.16 67.41 79.76 76.40 78.04 76s 140 15 536 4176
#10  PARFormer [9] TCSVT23 URL 75.33 74.21 79.33 85.82 82.09 732s 195 68 756 11634
#11  SequencePAR [19] arXiv23 URL 73.96 73.40 81.47 81.17 81.11 275s 466 413 1776 11263
#12  VTB-PLIP  [65] arXiv23 URL 73.72 71.50 79.97 80.57 79.99 26s 31 8 591 1614
#13  Rethink-PLIP [65] arXiv23 URL 57.91 62.78 78.27 70.81 74.35 44s 24 4 144 80
#14  PromptPAR [45] TCSVT24 URL 74.92 71.79 78.33 83.01 80.31 977s 8 3 1200 4435
#15  SSPNet  [63] PR24 URL 71.71 69.41 80.46 77.46 78.93 32s 26 7 100 1240
#16  HAP [57] NIPS24 URL 77.33 74.13 83.99 80.56 82.24 92s 86 17 327 2683
#17  MambaPAR [47] arXiv24 URL 72.28 70.57 77.99 81.70 79.50 18s 26 1 98 116
#18  MaHDFT [46] arXiv24 URL 75.54 72.27 80.28 82.19 80.59 158s 158 3978 431 187
              Ours - - 79.32 76.00 82.37 84.55 83.22 68s 107 39 429 2200
              Ours+E - - 87.66 84.78 89.03 89.38 89.07 361s 201 77 836 2312
5.2 Implementation Details
In our experiments, we use the VRWKV6-B version of the VRWKL [8] base model pre-trained on the ImageNet-1K [6] dataset as the visual encoder. This version consists of 12 block layers. We select SGD [12] as the optimizer, We leverage the warm-up strategy and increase the learning rate from 0 to the initial learning rate 0.008 linearly in the first 10 epochs, and decrease the learning rate by a factor of 0.1 when the number of iterations increases, we set the batch size to 16 and train for 60 epochs, the filtering threshold is set to 0.75. During both the training and inference stages, we first pad and resize the images to 256×128. Training images are augmented with random horizontal flipping with a probability of 0.5 and random cropping with a padding size of 10. Feature interaction is performed using the last layer of the transformer. Our model is implemented based on the PyTorch deep learning framework, and the experiments are conducted on a server equipped with an NVIDIA RTX 3090 GPU. For more details about our framework, please refer to our source code.
5.3 Comparison on Public Datasets
In this section, we conduct a detailed comparison between our model and existing pedestrian attribute recognition (PAR) algorithms on three benchmark datasets. Since the MARS-Attribute and DukeMTMC-VID-Attribute datasets contain only the RGB modality, we simulate the corresponding event data to ensure the completeness and fairness of the experiments.
$\bullet$ Results on EventPAR Dataset.  We collect and analyze public PAR methods from 2015 to 2024 on the EventPAR dataset as shown in Table 3, methods such as HAP [57], SequencePAR [19], and PARFormer [9] have demonstrated superior performance. Among them, the HAP method achieved outstanding results, with mA, Acc, Prec, Recall, and F1 scores of 77.33, 74.13, 83.99, 80.56, and 82.24, respectively. Our method achieves strong performance with mA, Acc, Prec, Recall, and F1 scores of 79.32, 76.00, 82.37, 84.55, and 83.22, respectively, outperforming existing approaches across most evaluation metrics. Furthermore, by incorporating the event auxiliary modality, our model further improves these scores to 87.66, 84.78, 89.03, 89.38, and 89.07, achieving the best overall results. These experimental findings on the EventPAR dataset not only confirm the effectiveness of our method in pedestrian attribute recognition tasks but also highlight the significant advantage of integrating event modality in enhancing model robustness under complex real-world conditions.
$\bullet$ Results on MARS-Attribute Dataset [61] As shown in Table 4, to further evaluate the generalization and effectiveness of our method, we conducted comprehensive experiments on another public dataset, comparing it with several state-of-the-art approaches. While our method did not consistently achieve the top performance across all metrics, it demonstrated strong competitiveness, achieving 73.74, 80.43, 84.63, and 82.18 in Accuracy, Precision, Recall, and F1, respectively. These results reflect a balanced performance with high accuracy and effective attribute capture.
$\bullet$ Results on DukeMTMC-VID-Attribute Dataset [34]. Our method achieves performance comparable to mainstream approaches on this dataset, with mA, Accuracy, Precision, Recall, and F1 of 68.19, 82.83, 77.30, 79.38, and 78.32, respectively. Although the overall performance is slightly lower than that of the VTB method, our approach still demonstrates solid generalization ability and stability. Notably, it achieves competitive results while maintaining relatively low computational complexity, validating the effectiveness and feasibility of our proposed method. These findings suggest that our method remains competitive in certain real-world scenarios and has the potential for further improvement through optimization of the model architecture and fusion strategy.
Table 4: Results on MARS-Attribute and DukeMTMC-VID-Attribute RGB-Event based PAR dataset.
Methods Backbone MARS-Attribute Dataset DukeMTMC-VID-Attribute Dataset
Accuracy Precision Recall F1 Accuracy Precision Recall F1
VTB [4] ViT-B/16 72.73 82.79 83.52 82.89 72.50 83.60 82.24 82.38
Zhou et al. [62] ConvNext 69.43 81.75 80.64 81.19 65.51 78.42 77.40 77.91
VTB-PLIP  [65] ResNet50 54.93 70.02 69.23 69.14 43.95 60.80 58.52 59.10
Rethink-PLIP [65] ResNet50 47.85 62.15 64.80 63.45 40.25 55.41 56.13 55.77
SSPNet  [63] ResNet50 65.68 77.09 79.74 78.39 67.53 78.54 79.78 79.15
MambaPAR [47] Vim 69.07 81.83 79.39 80.24 61.55 75.54 74.84 74.24
Ours RWKV-B 73.74 80.43 84.63 82.18 68.19 82.83 77.30 79.38
Table 5: Comparison of different input settings and filtering threshold is set to 0.95.
RGB Event mA Acc Prec Recall F1
1 0 79.32 76.00 82.37 84.55 83.22
0 5 87.10 84.49 89.03 89.07 88.91
1 5 87.66 84.78 89.03 89.38 89.07
3 5 85.95 81.87 86.80 86.95 86.56
5 5 85.67 82.20 87.13 86.86 86.72
Table 6: Comparison of different feature fusion methods.
Method mA Acc Prec Recall F1
Concat 87.63 84.60 88.80 89.40 88.96
Add 87.17 84.62 88.96 89.30 88.99
$1\times 1$ Conv 83.77 80.33 84.78 88.63 86.44
OTN-RWKV 87.66 84.78 89.03 89.38 89.07
Table 7: Comparison of different threshold setting in the feature filtering module.
Threshold mA Acc Prec Recall F1
0.60 87.23 84.04 88.62 88.91 88.62
0.65 87.55 84.16 88.62 89.09 88.70
0.70 87.68 83.93 88.32 89.11 88.71
0.75 87.66 84.78 89.03 89.38 89.07
0.80 87.70 84.13 88.75 88.92 88.69
0.85 87.72 84.74 88.99 89.40 89.05
0.95 86.55 82.43 87.32 87.44 87.05
Table 8: Ablation study of different backbones and framework modules on the EventPAR dataset.
No. Backbone Aggregation Fusion EventPAR Dataset
Max Mean Gnn Similarity Concat Add 1x1Conv OTN-RWKV mA Acc F1
01 RWKV 87.56 84.62 88.98
02 RWKV 87.50 84.68 89.01
03 RWKV 84.53 82.24 87.50
04 RWKV 87.59 84.80 89.10
05 RWKV 87.50 84.68 89.01
06 RWKV 86.92 84.45 88.88
07 RWKV 87.13 84.40 88.86
08 RWKV 87.38 84.82 89.10
09 ViT 84.05 82.41 87.60
10 ResNet50 83.85 82.45 87.64
11 RWKV 87.66 84.78 89.07
5.4 Ablation Study
In this section, we conduct a comprehensive ablation study on the core modules of EventPAR, including different input configurations, various backbone architectures, fusion strategies, event data aggregation methods, and the threshold settings in the similarity-based filtering mechanism.
$\bullet$ Analysis of different input settings. In this study, we varied the number of RGB inputs and conducted extensive experiments, as shown in Table 5. The first two rows represent RGB-only and Event-only inputs, both yielding strong results. RGB performed slightly worse due to added perturbations. When combining RGB and Event, performance improved, demonstrating their complementarity. However, excessive RGB inputs led to performance drops, likely due to redundant features and preprocessing noise. These factors increased the burden during feature extraction, introduced irrelevant cues, and weakened the ability to capture essential semantic information, ultimately degrading recognition performance.
$\bullet$ Analysis of different aggregation strategies of Event.  As shown in the first part of Table 8, we systematically compared four event data aggregation strategies: Max Pooling, Mean Pooling, a graph-based GNN, and our proposed Similarity-based approach. While Max and Mean pooling offer efficient but coarse feature representations, GNN captures structural dependencies with stronger expressive power. Our Similarity-based method adaptively fuses features based on semantic similarity between event frames, effectively retaining discriminative temporal information and filtering noise. n terms of mA 87.59, Accuracy 84.80, and F1 89.10, it outperformed all other strategies, hereby validating the effectiveness of our proposed similarity-based aggregation strategy.
$\bullet$ Analysis of different threshold setting in the similarity aggregation strategies  To further validate the impact of threshold settings in the similarity-based filtering strategy, we conducted experiments with varying similarity thresholds under identical conditions. The threshold controls the degree of fusion between event frames, preserving only those with sufficient semantic similarity to reduce redundant or noisy features. As shown in Table 7, a well-chosen threshold better preserves discriminative features and enhances performance, while extreme thresholds may impair fusion quality and degrade results.
$\bullet$ Analysis of different feature fusion methods.  We adopted two fusion strategies: holistic fusion and stepwise fusion. The corresponding results are shown in Table 6 and the second part of Table 8, respectively. We compared our proposed OTN-RWKV fusion strategy with three common methods: Concat, Add, and 1×1Conv. Although Concat increases feature dimensionality, it fails to capture inter-modal relationships. Add struggles with modality discrepancies, and 1×1Conv may lead to the loss of semantic details. In contrast, OTN-RWKV achieved superior performance in both fusion modes. It recorded mA of 87.66, ACC of 84.78, and F1 of 89.07 in holistic fusion, and mA of 87.38, ACC of 84.82, and F1 of 89.10 in stepwise fusion, outperforming all baseline methods. By finely integrating RGB and event modality features, the proposed strategy effectively captures their complementarity, significantly improves model robustness and overall performance, and fully validates its effectiveness.
$\bullet$ Analysis of different backbone.  To further evaluate the effectiveness of the proposed RWKV model, we conducted comparative experiments with ViT and ResNet-50 under identical settings. ViT captures global context, while ResNet-50 focuses on local features. In contrast, RWKV combines the strengths of RNNs and Transformers, achieving a balance between long-range dependency modeling and efficiency. Using consistent inputs and training protocols, our RWKV model outperformed both backbones with mA, Accuracy, and F1 of 87.66, 84.78, and 89.10, respectively, as shown in the third part of Table 8, highlighting its superiority in pedestrian attribute recognition.
Figure 6: Visualization of pedestrian attributes predicted by our proposed method. The red attributes indicate incorrect predictions, while green attributes represent ground truth.
5.5 Visualization
$\bullet$ Visualization of Predicted Results.  As shown in Fig. 6, we present a visualization of attribute prediction results using different input modalities: RGB only, Event only, and the fused bimodal input. In the figure, green text represents the ground truth attributes, while red text indicates incorrectly predicted ones. The results show that using a single modality, whether the RGB data with added noise or the Event data alone, cannot achieve comprehensive and accurate predictions. In contrast, when both modalities are combined, the model can better capture complementary information, which significantly improves the accuracy and robustness of attribute recognition.
$\bullet$ Visualization of Aggregation Strategy.  To provide an intuitive understanding of our proposed similarity-based aggregation strategy, we present a visual illustration in Fig. 7. In this figure, the colored regions represent event tokens that are retained and contribute to the final aggregation, while the black areas indicate tokens filtered out by the similarity-based selection process. As shown, the strategy effectively suppresses redundant or irrelevant information by adaptively retaining only semantically relevant tokens. This helps the model focus on discriminative patterns, thereby improving the overall recognition performance.
Figure 7: Illustration of our proposed similarity-based aggregation strategy for event token selection, where the black parts indicate the filtered-out tokens.
6 Conclusion and Future Works
In this paper, we addressed the limitations of existing pedestrian attribute recognition (PAR) methods, which are predominantly based on RGB frame cameras and suffer from challenges such as sensitivity to lighting conditions and motion blur. Moreover, current approaches focus primarily on external appearance and clothing attributes, neglecting the emotional dimensions of pedestrians. To overcome these constraints, we introduced a novel multi-modal RGB-Event attribute recognition task, leveraging the advantages of event cameras in low-light, high-speed, and low-power scenarios. Our key contribution is the creation of EventPAR, the first large-scale multi-modal PAR dataset, which includes 100K paired RGB-Event samples covering 50 attributes related to both appearance and six human emotions. By retraining and evaluating mainstream PAR models on EventPAR, we established a comprehensive benchmark, offering valuable data and algorithmic baselines for the community. Furthermore, we proposed a novel RWKV-based multi-modal PAR framework, incorporating an RWKV visual encoder and an asymmetric RWKV fusion module. Extensive experiments conducted on EventPAR and two simulated datasets demonstrated state-of-the-art performance, underscoring the effectiveness of our approach.
In our future works, we will focus on exploiting learning-based new event representations for the perception of event streams for pedestrian attribute recognition.